OpenStudy (anonymous):

The vertices of a parallelogram are A(x1, y1), B(x2, y2), C(x3, y3), and D(x4, y4). Which of the following must be true if parallelogram ABCD is proven to be a rectangle?

OpenStudy (anonymous):

(y4−y3x4−x3=y3−y2x3−x2) and (y4−y3x4−x3×y3−y2x3−x2)=-1 (y4−y3x4−x3=y2−y1x2−x1) and (y4−y3x4−x3×y2−y1x2−x1)=-1 (y4−y3x4−x3=y2−y1x2−x1) and (y4−y3x4−x3×y3−y2x3−x2)=-1 (y4−y3x4−x3=y3−y1x3−x1) and (y4−y3x4−x3×y2−y1x2−x1)=-1